OpenStudy (anonymous):

what value of a makes the equation true 24/40=9/a a.12 b.25 c.6 d.15

OpenStudy (anonymous):

can anyone help me break it down

OpenStudy (lieutenantgeneral):

simplify to 3/5 = 9/a = a = 15 cuz 9/15 divide both by 3 and 3/5